OpenStudy (abbycross167):

I will fan and give a medal!! What is 6 divided by -16? = -2.6?

OpenStudy (sweetnikkie072):

-0.375

OpenStudy (tommo_lcfc):

\[6/(-16) = -3/8 = -0.375\]

OpenStudy (ssddss02):

They are right

OpenStudy (abbycross167):

oh ok... I have another question that I need someone to check the steps would someone mind looking at it?

geerky42 (geerky42):

Sure

OpenStudy (sweetnikkie072):

go ahead

OpenStudy (abbycross167):

I will open a new question so I can give another medal

rebeccaxhawaii (rebeccaxhawaii):

it will be negative